Fermi National Laboratory
 

Operation Summary for June 17, 2011 - June 20, 2011

 

The Integrated Luminosity for the period from 6/13/11 to 6/20/11 was 78.44 pb-1 (inverse picobarns). NuMI reported receiving 5.64E18 protons on target during this same period.

Friday June 17

The day shift began with Operations monitoring the Tevatron (TEV) store 8820 (91.51E30) and the Recycler stash (323.4E10), with the Antiproton Source (Pbar) stacking (19.5E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, Operations terminated store 8820 at 9:05 AM. Operators turned off the TEV so that cryo system personnel could replace the TEV sector F4 wet engine flywheel. Operators prepared D0 for access. After all work and accesses were completed, Operations put the TEV into shot setup and established store 8824 at 3:25 PM with an average initial luminosity (AIL) of 380.35E30.

During the evening shift, operators reported a MI vacuum burst at 4:59 PM. At 5:13 PM, a Pbar person reported that an Accumulator RF station (ARF1) had a wandering output. Operators called in RF and Pbar personnel to investigate and they found a problem with the air-flow in the service building, which they adjusted. Operations disabled beam to T-978 at 6:05 PM.

Saturday June 18

The midnight shift began with Operations monitoring store 8824 (114.36E30) and the stash (235.4E10), with Pbar stacking (12.6E10), and with NuMI and MiniBooNE taking beam.

During the midnight shift, a Booster RF station (BRF6) tripped off and was reset for the first of several times during the shift. Operations established beam to T-978 at 4:21 AM. LRF1 tripped off for the first of five times in 20 minutes: it reset. Operations terminated store 8824 at 7:49 AM and put the TEV into shot setup.

During the day shift, Operations established store 8825 at 9 AM with and AIL of 413.18E30. Operators reported minor downtimes due to Booster and Linac RF trips.

During the evening shift, Operators reported at 5:15 PM that LRF5 tripped off for the first of four times in 20 minutes; it reset. Operations disabled beam to T-978 at 6:05 PM. Operators tuned the Linac 750 KeV line and Booster injection. Operators reported at 8:16 PM that a transfer of antiprotons from Pbar to the Recycler made it as far as MI and then was lost when a kicker misfired; operators reported that the next transfer was okay. BRF6 tripped off for the first of four times in five minutes; it reset. Operators reported at 11:05 PM that they were bypassing BRF6 and bringing BRF19 online in its place.

Sunday June 19

The midnight shift began with Operations monitoring store 8825 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations terminated store 8825 at 2:11 AM, put the TEV into shot setup, and established store 8826 at 3:26 AM with an AIL of 405.1E30. Operations established beam to T-978 at 4:48 AM. At 5:58 AM, operators reported that the Booster extraction kickers were causing serious beam position problems. Operators called in a Booster person to help investigate.

During the day shift, operators reported that a NuMI magnet power supply (E:V118) tripped off on a ground fault indication for the first of two times in 40 minutes; it reset. A Booster person reported at 8:28 AM that he replaced the trigger for the extraction kicker (MKS05). At 11:23 AM, a MI kicker power supply (I:KPS5S) tripped off on a Thyratron no-fire indication for the first of two times in 16 minutes. The first trip caused the loss of a transfer of antiprotons.

During the evening shift, Operations disabled beam to T-978 at 6:02 PM. NuMI tripped off at 7:45 PM when the target helium flow increased. Operators notified NuMI personnel and with their permission reset the beam permit; operators monitored the flow-rate. Operations terminated store 8826 at 7:53 PM and put the TEV into shot setup. Operations established store 8827 at 9:10 PM with an AIL of 390E30.

Monday June 20

The midnight shift began with Operations monitoring store 8827, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

During the midnight shift, Operations established beam to T-978 at 4:31 AM. Operators reported a nasty line of thunderstorms at 6 AM with lightning strikes that caused various system trips.

The Plans for Monday

The plans for today are to stack, stash, and store. The Run Coordinator will allow NuMI personnel to conduct target scans today in preparation for a target move tomorrow (6/21/11); operators will turn off beam to NuMI for the target move tomorrow. The experiment won’t resume taking normal beam until Wednesday morning (6/22/11) at 10 AM.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/20/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory